OpenStudy (anonymous):

If sin A = 2/3 where 0° < A < 90°, what is the value of sin 2A

OpenStudy (saifoo.khan):

41.81

myininaya (myininaya):

sin2A=2sinAcosA sinA=2/3 we need to know what cosA is sinA=opp/hyp=2/3 (opp)^2+(adj)^2=(hyp)^2 2^2+(adj)^2=3^2 (adj)^2=9-4 adj=sqrt{3} cosA=adj/hyp=sqrt{3}/3

myininaya (myininaya):

so sin2a=2*2/3*sqrt{3}/3=4sqrt{3}/9

OpenStudy (anonymous):

answer given is actually 4sqrt{5} /9

OpenStudy (anonymous):

dont know how to get it though

myininaya (myininaya):

oh strive do you see i subtract wrong above

myininaya (myininaya):

see that 9-4 i put 3 for some reason its suppose to be 5

myininaya (myininaya):

strive?

OpenStudy (anonymous):

ohhh gotcha